OpenStudy (anonymous):

please answer with solution 1.find an angle which is 1/2 its complement 2.find an angle which is twice its complement 3.two complementary angles have a difference of 18 degrees. find their measures. 4.two supplementary angles have a difference of 64degrees. find their measures. 5.find the measure of the larger of two supplementary angles if the measure of the larger is 4times the smaller.

OpenStudy (lgbasallote):

1. let x be one angle the other be y they are complements so x + y = 90 it says one is 1/2 the other so let's say x = 1/2 y \[\frac 12 y + y = 90\] if you combine like terms... \[\frac 12 y + \frac 22 y = 90\] \[\implies \frac 32 y = 90\] cross multiply... \[\implies 3y = 180\] divide both sides by 3... \\[\implies y = 60\] so the complement is 90 - 60 = 30 so the angles are 60 and 30 does that help? \